OpenStudy (anonymous):

How do you solve: e^–x + xe^–x = 0.123?

OpenStudy (anonymous):

factor e^-x

OpenStudy (anonymous):

\[e^{-x}(1+x)=0,123\] introduce ln

OpenStudy (anonymous):

when we use ln, does it become -x(1+x)=ln(0.123) ?

OpenStudy (paxpolaris):

it becomes : \[\large -x+\ln(x+1)=\ln(0.123)\] i don't know what to do from there.

OpenStudy (anonymous):

using log n then differentiate

OpenStudy (anonymous):

\[-x +\ln (1+x)=\ln 0,123\]

OpenStudy (phi):

you could solve this numerically.

OpenStudy (anonymous):

how

OpenStudy (shubhamsrg):

newton;s method for approximation might help maybe,,but its quite tedious..

OpenStudy (anonymous):

woaaah?

OpenStudy (phi):

newton's method works fine. But wolfram is easier, unless you have to solve this by hand

OpenStudy (phi):

If you wanted to do it with a program, here is an example. The difficult part is knowing where to start the search for the 2 roots, so plotting the function is helpful.